OpenStudy (anonymous):

(x+3)^3-(3x-2)^2=x^3+7

OpenStudy (amoodarya):

what grade are you in ? you know derivation ?

OpenStudy (anonymous):

im in college and no sorry i don't know what that is

OpenStudy (anonymous):

Move the RHS to the LHS, multiply out the products and simplify.\[16 + 39 x = 0 \]\[x=-\frac{16}{39} \]

OpenStudy (anonymous):

ok i have another question for you

OpenStudy (anonymous):

Go ahead.

OpenStudy (anonymous):

10/y^2-16 - 3/y+4 =4/y-4

OpenStudy (anonymous):

how would i do that. i know have to multiply by (y-4) and (y+4) but after that i get lost

OpenStudy (anonymous):

no its \[10\y^2-16\] - \[3\y+4\]=\[4\y-4\]

OpenStudy (anonymous):

Sorry, I have run out of time. Have to sign off from this site.